OpenStudy (acxbox22):

PLease help....medal.... Suppose you are the treasurer of the drama club. the cost of scripts for the spring musical is $254.the cost of props and costumes is $400. you must also pay $1.20 per ticket to the play's director.you charge $4.00 per ticket and you also expect to make $150 on refreshments. How many tickets do you need to sell to break even?

OpenStudy (acxbox22):

@jim_thompson5910 , @SolomonZelman , @Hero

OpenStudy (acxbox22):

@ranga

OpenStudy (acxbox22):

@austinL

OpenStudy (mathmale):

"Break even," in this situation, means that Costs = Revenues. You'll need to come up with linear equations for Costs: C(x), and Revenues: R(x). Look at the problem statement. What components (parts) does the Cost function have?

OpenStudy (mathmale):

It costs money to make money! What are the costs involved in this situation?

OpenStudy (acxbox22):

400+254+1.2 per ticket

OpenStudy (acxbox22):

right @mathmale ...

OpenStudy (ranga):

Assume you need to sell "x" tickets to break even. What is the total cost if "x" tickets are sold? What is the total revenue if "x" tickets are sold? equate the two and solve for x.

OpenStudy (acxbox22):

654+1.2x=4x+150 504=2.8x x=180 ??

OpenStudy (mathmale):

To answer your question: 400+254+1.2 per ticket is correct. But since we need to find out how many tickets need to be solved to "break even," write C(x) as C(x)=400+254+1.2x. Now, what about an expression for Revenue? I see you've got that. I'm not doing the actual math myself. But what you have written appears to be completely correct. What does that 180 signify?

OpenStudy (mathmale):

Reread the original problem statement for information on that.

OpenStudy (acxbox22):

180 signfies the tickets needed to equal revenue and expenditure

OpenStudy (acxbox22):

thanks mathmale, i was just doing the end of the year review for state test for algebra and i remember how to do this now